# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v4

@@ Abstract.
We establish a quantitative relationship between mixed cohomology classes and the geometric complexity of cohomologically calibrated metric connections with totally skew torsion on product manifolds. We show that the dimension of the off-diagonal curvature subspace of a connection $\nabla^{C}$ is bounded below by the sum of tensor ranks of the mixed Künneth components of its calibration class, minus the dimension of the space of harmonic parallel forms on the first factor. The bound depends only on the mixed class $\omega]_{\mathrm{mixed}}\in H^{3}(M;\mathbb{R})$ and on the space of $g_{1}$-parallel harmonic $2$-forms on $M_{1}$; it is therefore independent of the factor metric $g_{2}$ and of any deformation of $g_{1}$ that preserves parallel $2$-forms, hence it is metric-invariant modulo parallel-form preserving deformations. This provides a computational criterion for geometric complexity and quantifies the interaction between topology and curvature, yielding a quantified version of “forced irreducibility” via the dimension of $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$.
@@ 1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological quantity, the mixed tensor rank of the cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that it furnishes, for any metric connection $\nabla^{\mathcal{C}}$, a quantified lower bound for the dimension of the subspace of the holonomy algebra generated by off-diagonal curvature operators with respect to the splitting $V_{1}\oplus V_{2}$. The key point is that, although the definition of “off-diagonal” uses $g$, the lower bound is independent of the factor metric $g_{2}$ and of any deformation of $g_{1}$ that preserves the space of parallel 2-forms, and is explicitly corrected by the dimension of $g_{1}$-parallel harmonic 2-forms on the first factor. Each independent mixed Künneth component in $\omega]$ guarantees a linearly independent off-diagonal curvature direction, modulo the parallel subspace.
@@ 2. Cohomologically Calibrated Metric Connections with Skew Torsion
A metric connection $\nabla^{C}$ with non-parallel ($\nabla^{LC}T\neq 0$) and totally skew torsion $T$, where the cohomology class $\omega]$ is non trivial and defined via the harmonic projection of the torsion three-form, ($T^{\flat}=\omega_{h}+d\alpha+\delta\beta$, so $\omega]:=[\omega_{h}]$), is called cohomologically calibrated. The class is mixed if it lies outside the natural image of $H^{3}(M_{1})\oplus H^{3}(M_{2})$ under the Künneth isomorphism:
@@ 3. Off-Diagonal Curvature and Holonomy Complexity
The off-diagonal holonomy subspace at $p\in M$ is
$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\operatorname{Span}_{\mathbb{R}}\{\,R_{\mathrm{off}}(X,Y)Z\;:\;X,Y,Z\in T_{p}M\,\}\ \subseteq\ \mathfrak{hol}_{p}(\nabla^{C}).$ (3.2)
Clearly $\dim\mathfrak{hol}_{p}(\nabla^{C})\geq\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$. Because $\nabla^{C}$ is metric, its holonomy algebra sits in $\mathfrak{so}(T_{p}M)$. If the action were reducible with respect to $V_{1}\oplus V_{2}$, all holonomy endomorphisms would be block-diagonal, forcing $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\{0\}$. Hence nontrivial off-diagonal holonomy implies irreducibility.
@@ 5. Main Result and Proof
Let $(M_{1},g_{1})$ be a closed oriented smooth Riemannian manifold and $\mathcal{H}^{2}(M_{1})$ the space of harmonic $2$-forms. Fix an $L^{2}$-orthonormal basis $\{\alpha_{i}\}_{i=1}^{r_{2,1}}$ of a subspace of $\mathcal{H}^{2}(M_{1})$ such that the first $s:=\dim\mathcal{P}_{2}$ elements span the parallel subspace $\mathcal{P}_{2}$. Then the set
$\mathcal{U}^{\prime}:=\Bigl\{p_{1}\in M_{1}\;\Big|\;\text{rank}\!\bigl[(\nabla^{g_{1}}\alpha_{i})(p_{1})\bigr]_{i=s+1,\dots,r_{2,1}}=r_{2,1}-s\Bigr\}$
@@
The condition $\text{rank}\mathbf{A}(p_{1})=r_{2,1}-s$ is equivalent to the non-vanishing of at least one $(r_{2,1}-s)\times(r_{2,1}-s)$ minor of $\mathbf{A}(p_{1})$. Each minor is a smooth function on $B$; hence its non-zero locus is open in $B$. Covering $M_{1}$ by such balls we conclude that $\mathcal{U}^{\prime}$ is open. Suppose, by contradiction, that $\mathcal{U}^{\prime}$ is not dense. Then there exists a non-empty open set $V\subset M_{1}$ on which
$\text{rank}\mathbf{A}(p_{1})<r_{2,1}-s\qquad\forall\,p_{1}\in V.$
Thus, for every $p_{1}\in V$, there exists a non-zero vector $\lambda=(\lambda^{s+1},\dots,\lambda^{r_{2,1}})\in\mathbb{R}^{r_{2,1}-s}$ (depending on $p_{1}$) such that
$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(\nabla^{g_{1}}\alpha_{i})(p_{1})=0.$
The smooth section
$\eta:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}\alpha_{i}\in\mathcal{H}^{2}(M_{1})$
then satisfies $\nabla^{g_{1}}\eta\equiv 0$ on $V$. By unique continuation for harmonic forms ([1]), a harmonic form parallel on an open set of any connected component of $M_{1}$ is globally parallel on that component (By Greene–Jacobowitz [4], we approximate $g$ in $C^{\infty}$ by real-analytic metrics $g_{k}$; for each such metric the Hodge-Laplace operator has analytic coefficients, so Aronszajn’s unique-continuation theorem applies to harmonic $2$-forms). Hence, $\eta$ is parallel on the connected component containing $V$, which implies $\eta\in\mathcal{P}_{2}$. This contradicts the linear independence of $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ modulo $\mathcal{P}_{2}$. Therefore $\mathcal{U}^{\prime}$ must be dense in $M_{1}$.
@@
Let $M=M_{1}\times M_{2}$ be compact and oriented, $g=g_{1}\oplus g_{2}$ any product metric, and $\nabla^{C}$ any cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]:=[\omega_{h}]\in H^{3}(M;\mathbb{R})$ is a mixed class.
Then for every $p\in M$,
$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\geq\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\!\bigl(\mathcal{P}_{2}\otimes H^{1}(M_{2})\bigr),$
where $\mathcal{P}_{2}\subset H^{2}(M_{1};\mathbb{R})$ denotes the space of harmonic $2$-forms on $M_{1}$ that are parallel with respect to $g_{1}$.
The bound depends only on $\omega]_{\mathrm{mixed}}$ and on the space of $g_{1}$-parallel harmonic $2$-forms; it is therefore invariant under any deformation of $g_{2}$ and under $g_{1}$-deformations that preserve parallel $2$-forms, hence metric-invariant modulo parallel-form preserving deformations.
@@
with $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$ harmonic and chosen so that the number of terms is minimal. Similarly, write
@@
with $\tilde{\alpha}_{j}\in\mathcal{H}^{1}(M_{1})$ and $\tilde{\beta}_{j}\in\mathcal{H}^{2}(M_{2})$. The total number of simple mixed tensors equals $r:=r_{2,1}+r_{1,2}=\mathrm{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)$.
Define the parallel subspace as
$\mathcal{P}_{2}:=\bigl\{\alpha\in\mathcal{H}^{2}(M_{1})\mid\nabla^{g_{1}}\alpha=0\bigr\}\subset H^{2}(M_{1};\mathbb{R})$ (5.3)
and set
$r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)-\dim\!\big(\mathcal{P}_{2}\otimes H^{1}(M_{2})\big).$ (5.4)
We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
@@
Consider the curvature formula (2.5). The mixed contribution of $T=\omega_{h}+\text{(coexact/exact)}$ to $R^{C}$ has two sources. First, the covariant derivative term $\nabla^{LC}T$ produces operators that, when $X,Y$ lie in $V_{1}$ and $Z$ in $V_{2}$ (or vice versa), insert a mixed $3$-form and hence map between $V_{1}$ and $V_{2}$. Second, the quadratic term $Q_{T}$ contains contractions of two copies of $T$ which, for mixed arguments, again produce off-diagonal operators. Pointwise at $p$, for each simple mixed component $\alpha_{i}\wedge\beta_{i}$ one can choose $X,Y\in V_{1}$ so that $\alpha_{i}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\beta_{i}(Z)\neq 0$. Substituting into (2.5) yields an endomorphism with a nontrivial $V_{2}\to V_{1}$ component; dually, for each $\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j}$ one chooses $X\in V_{1}$, $Y,Z\in V_{2}$ to obtain a nontrivial $V_{1}\to V_{2}$ component.
@@
To prove linear independence, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor-rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists an open-dense subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X,Y\in T_{p_{1}}M_{1}$ and $Z\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X,Y)|_{p}=\delta_{ik}$ and $\beta_{i}(Z)|_{p}=\delta_{il}$, obtaining $r$ linearly independent endomorphisms $R_{kl}^{\mathrm{off}}(p)\in\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$. This produces at least $r$ linearly independent elements in $\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})$, since curvature endomorphisms generate the holonomy algebra by Ambrose–Singer [2].
For each simple summand $\alpha_{i}\otimes\beta_{i}]$ of the minimal tensor-rank decomposition, choose harmonic representatives $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$. By unique continuation for harmonic forms, their pointwise evaluations are linearly independent on an open-dense subset $\mathcal{U}\subset M$. For every $p=(p_{1},p_{2})\in\mathcal{U}$ the off-diagonal component of the linear part of the curvature,
$R^{\mathrm{off}}(X,Y)Z=\tfrac{1}{2}\big(\nabla^{LC}_{p}T\big)^{\mathrm{off}}(X,Y,Z),$
evaluated on vectors $X,Y\in T_{p_{1}}M_{1}$ and $Z\in T_{p_{2}}M_{2}$, defines a linear map

### v9

@@ Abstract.
We establish a quantitative relationship between mixed cohomology classes and the geometric complexity of metric connections with totally skew torsion on product manifolds. For any such connection whose non‑parallel torsion 33‑form represents a mixed de Rham class via its harmonic projection, we prove that the dimension of the off‑diagonal holonomy subspace satisfies the lower bound
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the linear obstruction maps $\Psi_{p},\widetilde{\Psi}_{p}$. These kernels are defined intrinsically as the intersections
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\quad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (respectively $\mathcal{V}_{1,2}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2})$) denotes the linear span of the harmonic representatives of the $(2,1)$- (respectively $(1,2)$-) Künneth component of $\omega]$. The subspaces $\mathcal{V}_{2,1}$ and $\mathcal{V}_{1,2}$ are independent of the choice of a minimal harmonic decomposition. Consequently, the integer $\dim\mathcal{K}$ depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics; hence the bound is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata. This yields a computable, metric‑invariant obstruction to reducible holonomy and quantifies the topological “mixing strength” that forces irreducibility across the product splitting.
@@ 1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
More precisely, we prove
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K},$
where $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$ is the direct sum of the kernels of the natural linear obstruction maps. To define these kernels, let
$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp. $\{\widetilde{\alpha}_{j}\otimes\widetilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp. $(1,2)$-) Künneth component of $\omega]$. Because any two minimal harmonic decompositions span the same subspaces $U\subset\mathcal{H}^{2}(M_{1})$ and $V\subset\mathcal{H}^{1}(M_{2})$, the spaces $\mathcal{V}_{2,1}=U\otimes V$ and $\mathcal{V}_{1,2}$ are well‑defined, i.e., independent of the chosen decomposition. Then
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
The integer $\dim\mathcal{K}$ measures precisely the part of the mixed class that is absorbed by the parallel‑form strata and therefore does not contribute to off‑diagonal curvature. The quantity $r^{\sharp}$ is metric‑invariant under deformations that preserve the parallel‑form strata: it depends only on $\omega]_{\mathrm{mixed}}$ and on the subspaces $\mathcal{P}_{k}(M_{i})$. Consequently, each independent mixed Künneth tensor guarantees a linearly independent off‑diagonal holonomy direction, modulo the parallel forms captured by $\mathcal{K}$.
This yields a computable, purely topological criterion for forced irreducibility of the holonomy representation across the product splitting: irreducibility is forced whenever $r^{\sharp}>0$, i.e. whenever $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})>\dim\mathcal{K}$.
@@ 2. Cohomologically Calibrated Metric Connections with Skew Torsion
A metric connection $\nabla^{C}$ with non-parallel ($\nabla^{LC}T\neq 0$) and totally skew torsion $T$, where the torsion 3-form represents a non trivial mixed cohomology class $\omega]$ via its harmonic projection ($T^{\flat}=\omega_{h}+d\alpha+\delta\beta$, i.e. $\omega]:=[\omega_{h}]$), is called cohomologically calibrated. The class is mixed if it lies outside the natural image of $H^{3}(M_{1})\oplus H^{3}(M_{2})$ under the Künneth isomorphism:
@@ 3. Off-Diagonal Curvature and Holonomy Complexity
Let $(M,g)$ be a Riemannian manifold and $\nabla^{C}$ a metric connection on $TM$ with totally skew–symmetric torsion. For a point $p\in M$, we define the off–diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ as the Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ generated by all off–diagonal components of the curvature tensor and its covariant derivatives at $p$, namely
$\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C}):=\mathrm{Lie}\,\Big\langle\big(\nabla^{k}R^{C}\big)_{p}(X_{1},\ldots,X_{k};Y,Z)^{\mathrm{off}}\;\Big|\;k\geq 0,\;X_{i},Y,Z\in T_{p}M\Big\rangle,$
where $(\cdot)^{\mathrm{off}}$ denotes the off–diagonal projection with respect to the product splitting $T_{p}M=V_{1}\oplus V_{2}$.
By the Ambrose–Singer theorem [2], $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ depends only on the global holonomy algebra of $\nabla^{C}$ and is therefore independent of the choice of local generators and of the pointwise rank of the curvature tensor. In particular, at each point $p\in M$ one has, clearly, $\dim\mathfrak{hol}_{p}(\nabla^{C})\geq\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$. Because $\nabla^{C}$ is metric, its holonomy algebra sits in $\mathfrak{so}(T_{p}M)$. If the action were reducible with respect to $V_{1}\oplus V_{2}$, all holonomy endomorphisms would be block-diagonal, forcing $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\{0\}$. Hence nontrivial off-diagonal holonomy implies irreducibility.
@@ 5. Main Result and Proof
Let $(M_{1},g_{1})$ be a compact and oriented smooth Riemannian manifold and $\mathcal{H}^{2}(M_{1})$ the space of harmonic $2$-forms. Fix an $L^{2}$-orthonormal basis $\{\alpha_{i}\}_{i=1}^{r_{2,1}}$ of a subspace of $\mathcal{H}^{2}(M_{1})$ such that the first $s:=\dim\mathcal{P}_{2}$ elements span the parallel subspace $\mathcal{P}_{2}$. Then the set
$\mathcal{U}^{\prime}:=\Bigl\{p_{1}\in M_{1}\;\Big|\;\operatorname{rank}\!\bigl[(\nabla^{g_{1}}\alpha_{i})(p_{1})\bigr]_{i=s+1,\dots,r_{2,1}}=r_{2,1}-s\Bigr\}$
@@
The condition $\operatorname{rank}\mathbf{A}(p_{1})=r_{2,1}-s$ is equivalent to the non‑vanishing of at least one $(r_{2,1}-s)\times(r_{2,1}-s)$ minor of $\mathbf{A}(p_{1})$. Each minor is a smooth function on $B$; hence its non‑zero locus is open in $B$. Covering $M_{1}$ by such balls we conclude that $\mathcal{U}^{\prime}$ is open.
Now suppose, for a contradiction, that $\mathcal{U}^{\prime}$ is not dense. Then there exists a non‑empty open set $V\subset M_{1}$ on which
$\operatorname{rank}\mathbf{A}(p_{1})<r_{2,1}-s\qquad\forall\,p_{1}\in V.$
Since the rank is lower‑semicontinuous, after possibly shrinking $V$ we may assume that the rank is constant, say $\rho<r_{2,1}-s$, on $V$. Consequently, there exists a smooth map $\lambda:V\to\mathbb{R}^{r_{2,1}-s}\setminus\{0\}$ such that
$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{1})(\nabla^{g_{1}}\alpha_{i})(p_{1})=0\qquad\forall p_{1}\in V.$ (5.1)
Now we employ analytic approximation. By the theorem of Greene–Jacobowitz [4], we can approximate the metric $g_{1}$ in the $C^{\infty}$ topology by a sequence of real‑analytic metrics $\{g_{1}^{(k)}\}$. For each $k$, denote by $\alpha_{i}^{(k)}$ the $g_{1}^{(k)}$-harmonic representative of the cohomology class $\alpha_{i}]$. The forms $\alpha_{i}^{(k)}$ converge to $\alpha_{i}$ in $C^{\infty}$ as $k\to\infty$.
Because $\lambda$ is smooth and non‑zero on $V$, we may choose a point $p_{0}\in V$ and a subsequence (still denoted by $k$) such that $\lambda(p_{0})\neq 0$ and
$\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})(\nabla^{g_{1}^{(k)}}\alpha_{i}^{(k)})(p_{0})=0\qquad\text{for all }k.$ (5.2)
Indeed, (5.1) passes to the limit after extracting a suitable subsequence, since all objects depend smoothly on the metric.
For each fixed $k$, the metric $g_{1}^{(k)}$ is real‑analytic. Define
$\eta^{(k)}:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}^{(k)}\in\mathcal{H}^{2}(M_{1},g_{1}^{(k)}).$
Equation (5.2) states that $\nabla^{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$. Because $g_{1}^{(k)}$ is analytic, the coefficients of the Hodge–Laplacian are analytic, and Aronszajn’s unique-continuation theorem [1] applies. Since $\eta^{(k)}$ is harmonic and $g_{1}^{(k)}$ is real-analytic, $\eta^{(k)}$ is also real-analytic. The condition $\nabla_{g_{1}^{(k)}}\eta^{(k)}(p_{0})=0$ implies that $\eta^{(k)}$ is parallel in a neighborhood of $p_{0}$. By the unique-continuation principle for parallel forms on real-analytic manifolds (a consequence of Aronszajn’s theorem applied to the analytic coefficients), $\eta^{(k)}$ must be parallel on the entire connected component.
Passing to the limit $k\to\infty$, we obtain that
$\eta:=\sum_{i=s+1}^{r_{2,1}}\lambda^{i}(p_{0})\,\alpha_{i}$
is parallel with respect to the original metric $g_{1}$ on the component containing $p_{0}$. Thus $\eta\in\mathcal{P}_{2}(M_{1})$, contradicting the linear independence of $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ modulo $\mathcal{P}_{2}(M_{1})$.
Therefore $\mathcal{U}^{\prime}$ must be dense in $M_{1}$.
The same reasoning applies verbatim to $\mathcal{H}^{k}(M_{i})$ for any $k$ and for both factors $M_{1}$ and $M_{2}$. ∎
@@
Let $M=M_{1}\times M_{2}$ be a compact and oriented product manifold with a product metric $g=g_{1}\oplus g_{2}$, and let $\nabla^{C}$ be a cohomologically calibrated metric connection with totally skew torsion $T$ such that $\omega]=[\omega_{h}]\in H^{3}(M;\mathbb{R})$ is a mixed class.
Let
$\mathcal{V}_{2,1}:=\operatorname{span}\{\alpha_{i}\otimes\beta_{i}\}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2}),\qquad\mathcal{V}_{1,2}:=\operatorname{span}\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}\subset\mathcal{H}^{1}(M_{1})\otimes\mathcal{H}^{2}(M_{2}),$
where $\{\alpha_{i}\otimes\beta_{i}\}$ (resp. $\{\tilde{\alpha}_{j}\otimes\tilde{\beta}_{j}\}$) is any minimal harmonic decomposition of the $(2,1)$- (resp. $(1,2)$-) Künneth component of $\omega]$. Define
$\ker\Psi_{p}:=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}:=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
and set $\mathcal{K}:=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}$. Then, for every $p\in M$,
$\dim\bigl(\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\bigr)\geq r^{\sharp}:=\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}.$
The bound depends only on the mixed class $\omega]_{\mathrm{mixed}}$ and on the parallel‑form strata $\mathcal{P}_{k}(M_{i})$; hence it is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata.
@@
with $\alpha_{i}\in\mathcal{H}^{2}(M_{1})$ and $\beta_{i}\in\mathcal{H}^{1}(M_{2})$ harmonic and chosen so that the number of terms is minimal. Similarly,
@@
with $\tilde{\alpha}_{j}\in\mathcal{H}^{1}(M_{1})$ and $\tilde{\beta}_{j}\in\mathcal{H}^{2}(M_{2})$. The total number of simple mixed tensors equals $r:=r_{2,1}+r_{1,2}=\operatorname{rank}_{\mathbb{R}}\big([\omega]_{\mathrm{mixed}}\big)$.
Set $r^{\sharp}:=r-\dim\mathcal{K}$. We will prove $\dim\mathfrak{hol}_{p}^{\operatorname{off}}(\nabla^{C})\geq r^{\sharp}$.
@@
Consider the curvature formula (2.5). The mixed contribution of $T=\omega_{h}+\text{(coexact/exact)}$ to $R^{C}$ has two sources. First, the covariant derivative term $\nabla^{LC}T$ produces operators that, when $X,Y$ lie in $V_{1}$ and $Z$ in $V_{2}$ (or vice versa), insert a mixed $3$-form and hence map between $V_{1}$ and $V_{2}$. Second, the quadratic term $Q_{T}$ contains contractions of two copies of $T$ which, for mixed arguments, again produce off‑diagonal operators. Pointwise at $p$, for each simple mixed component $\alpha_{i}\wedge\beta_{i}$ one can choose $X,Y\in V_{1}$ so that $\alpha_{i}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\beta_{i}(Z)\neq 0$. Substituting into (2.5) yields an endomorphism with a nontrivial $V_{2}\to V_{1}$ component; dually, for each $\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j}$ one chooses $X\in V_{1}$, $Y,Z\in V_{2}$ to obtain a nontrivial $V_{1}\to V_{2}$ component.
@@
To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists an open‑dense subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
Define the linear map